Cannon Co had Gross Profit Margin equal to 32.1% for the period ended 12/31/2008. Cannon had Sales of$3,074,000, what was Cannon’s COGS?
Answer #1

Gross Profit %= Gross Profit/Sales

Gross Profit=(3,074,000*32.1%)=986754

Gross Profit=Sales-Cost of goods sold

Cost of goods sold=3,074,000-986754

=$2087246